OpenStudy (anonymous):

can some one check my answer Please! answer 7/2 or 1/10

OpenStudy (anonymous):

(5-6)^2+2/5

OpenStudy (anonymous):

7/5

OpenStudy (radar):

\[(5-6)^{2}+2/5=(-1)^{2}+2/5=1+2/5=5/5 + 2/5 = 7/5\] \[(5-6)^{2}+2/5=(25-60+36) +2/5=(61-60) + 2/5=1 + 2/5=5/5 + 2/5= 7/5\